k^3-6k^2-k+6 = 0
Enter any solution as an integer or reduced fraction (not a mixed number or decimal), using a comma if there is more than one solution. If there is no solution, click No solution. Suggestion: Check any solution(s) in the original equation before submitting answer.

Found 2 solutions by mananth, MathLover1:
Answer by mananth(16946)


k^3-6k^2-k+6 = 0
k^2(k-6)-1(k-6)=0
(k^2-1)(k-6)=0
(k+1)(k-1)(k-6)=0

k= 1, -1,6



Answer by MathLover1(20849)